P lb Q lb .5 0.5L F lb COLLAPSE IMAGES Q=192b P#112b F:53 lb h=7ft L=13ft α=27 deg β-80 deg The reaction at A. lb ENTER 3 tries remaining. 1 point(s) possible The vertical reaction at E lb ENTER 3 tries remaining. 1 point(s) possible The horizontal reaction at E NIR
Answer #1

023 3.S6 13 6.5 53 13 터. + AR Eq+0. g& AR 19-48-17-073 オ0.884:R=-151.625/--② 彡H 6.163 AR + 1197-512ニ 344-5ナ136-14+119a . 512.

6-163 AR 1613:8 ey = 250.464-0.474x 26,353 一一3:82.095ル ,